the sum of the age of father and son is now 48 , what was the sum of their age ten years ago?
Answer by stanbon(75887)

the sum of the age of father and son is now 48 , what was the sum of their age ten years ago?
------
Equation:
f + s = 48
-----
f-10 + s-10 = 48-20
----
Answer: 28 yrs
==========
Cheers,
Stan H.
==========